Hello, can you please help me solve for x? x-2/x+6 = x/2. I came up with x=0.
Thank you!

Answer by stanbon(75887)

x-2/x+6 = x/2
Cross-multiply to get:
2(x-2) = x(x+6)
2x-4 = x^2+6x
x^2+4x+4 = 0
Factor to get:
(x+2)(x+2)=0
x+2=0
x = -2
===================
Cheers,
Stan H.